OpenStudy (anonymous):

Find the equation of the line through points (-8,5) and (-8,8) Help!

OpenStudy (anonymous):

first we need to find the slope. Do you know how to do that ?

OpenStudy (anonymous):

no :/

OpenStudy (jdoe0001):

\(\bf slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

there is a formula : slope(m) = (y2 - y1) / (x2 - x1) and your points : (-8,5) x1 = -8 and y1 = 5 (-8,8) x2 = -8 and y2 = 8 now just sub our stuff into the formula... slope(m) = (8 - 5) / (-8 - (-8) slope(m) = 3/0 This is an undefined slope. This means that the line is vertical, therefore we only need the x coordinate. What is the x coordinate in both sets of points ?

OpenStudy (anonymous):

the x coordinate is -8?

OpenStudy (anonymous):

correct. So your equation is x = -8

OpenStudy (anonymous):

So x= -8 would be the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wow. ok, thank you!

OpenStudy (anonymous):

no problem :)